# Neural Configuration Distance Function for Continuum Robot Control

### v1

@@ Abstract
This paper presents a novel method for modeling the shape of a continuum robot as a Neural Configuration Euclidean Distance Function (N-CEDF). By learning separate distance fields for each link and combining them through the kinematics chain, the learned N-CEDF provides an accurate and computationally efficient representation of the robot’s shape. The key advantage of a distance function representation of a continuum robot is that it enables efficient collision checking for motion planning in dynamic and cluttered environments, even with point-cloud observations. We integrate the N-CEDF into a Model Predictive Path Integral (MPPI) controller to generate safe trajectories. The proposed approach is validated for continuum robots with various links in several simulated environments with static and dynamic obstacles.
@@ I Introduction
We demonstrate that the advantages of the N-CEDF representation by integrating it into a Model Predictive Path Integral (MPPI) controller [20]. This enables safe and efficient motion planning for continuum robots relying on point cloud observations of dynamic and cluttered environments.
@@ III Problem Formulation
Figure 1: (a) A single continuum robot link with parameters: cable lengths $l_{1},l_{2},l_{3}$, backbone length $L$, bending angle $\theta$, and bending plane angle $\varphi$. (b) A 4-link continuum robot with a specific configuration in a dynamic environment with spherical obstacles and an end-effector goal (blue star).
A 3D continuum robot is modeled as a series of $M$ deformable links [4, 10, 5]. Each link consists of a flexible, inextensible backbone of length $L_{i}$ driven by three cables equispaced at intervals of $\frac{2\pi}{3}$ radians around the backbone (Fig. 1(a)). By controlling the cable lengths $l_{i,j}$, where $i\in\{1,\dots,M\}$ and $j\in\{1,2,3\}$, the spatial configuration of each link and the entire robot body can be actively manipulated. We define $l_{\min}$ and $l_{\max}$ as the lower and upper bounds of the cable length, respectively, such that $l_{\min}\leq l_{i,j}\leq l_{\max}$ holds. With the constant curvature model, the lengths of the three drive cables are constrained by the relationship $\frac{1}{3}\sum_{j=1}^{3}l_{i,j}=L_{i}$ for all $i$.
According to the PCC model, the bending of each link can be described by three curve parameters: the radius $\rho_{i}$ of the circular arc of the backbone, the bending angle $\theta_{i}\in[0,\pi]$, and the bending plane angle $\varphi_{i}\in[-\pi,\pi)$. These parameters can be derived from the cable lengths as follows [9, 5]:
@@
Let $\mathbf{x}^{k}=[l_{1,1}^{k},l_{1,2}^{k},l_{1,3}^{k},\dots,l_{M,1}^{k},l_{M,2}^{k},l_{M,3}^{k}]^{\top}\in\mathbb{R}^{3M}$ denote the vector containing the lengths of all drive cables at time step $k$. The discrete-time robot dynamics are:
@@
where $\mathbf{u}^{k}\in\mathbb{R}^{3M}$ is the control input, and $\tau^{k}$ is the sampling time. The relation between $\mathbf{q}^{k}$ and $\mathbf{x}^{k}$ is given in (1).
We consider an environment containing both static and dynamic obstacles. Let ${\cal O}^{k}\subset\mathbb{R}^{3}$ denote the closed obstacle set at time step $k$, and ${\cal F}^{k}=\mathbb{R}^{3}\setminus{\cal O}^{k}$ represent the free space, which is an open set.
@@ IV-B Data Preparation and Loss Function
To train $\hat{\Gamma}_{i}(\mathbf{p},\mathbf{q}_{i};\boldsymbol{\theta}_{i})$, we generate the following dataset. First, we uniformly sample a set of workspace points $\mathcal{P}_{w}=\{\mathbf{p}_{1}^{w},\dots,\mathbf{p}_{N_{w}}^{w}\}$ around the $i$-th link. We also uniformly sample a set of link configurations ${\cal Q}=\{\mathbf{q}_{i}^{1},\dots,\mathbf{q}_{i}^{N}\}$, where each $\mathbf{q}_{i}^{j}$ is sampled from the valid configuration space defined by the cable length limits. Next, given the configuration $\mathbf{q}_{i}^{j}$, we uniformly sample a set of points $\mathcal{P}_{s}^{j}(\mathbf{q}_{i}^{j})=\{\mathbf{p}_{1}^{s},\dots,\mathbf{p}_{N_{s}}^{s}\}$ on the surface of the link $\mathcal{B}_{i}(\mathbf{q}_{i}^{j})$. The Euclidean distance from each workspace point $\mathbf{p}_{m}^{w}\in\mathcal{P}_{w}$ to the link with configuration $\mathbf{q}_{i}^{j}$ is computed by:
@@ V-B Cost Function Design
The cost function is composed of three terms: goal-reaching cost $C_{\text{goal}}(\mathbf{Q}^{j})$, collision avoidance cost $C_{\text{coll}}(\mathbf{Q}^{j})$, and state constraint violation cost $C_{\text{state}}(\mathbf{Q}^{j})$. The goal-reaching cost penalizes the distance between the end-effector and the goal. The collision avoidance cost penalizes the robot being too close to obstacles, utilizing the learned N-CEDF model and kinematics chains. The state constraint violation cost penalizes the controlled cable lengths exceed their allowable limits. The individual terms are defined as:
@@
where $\|\cdot\|_{F}$ is the Frobenius norm, $w_{\text{goal}}$, $w_{\text{coll}}$, and $w_{\text{state}}$ are tunable weights, $l_{i,m}(\mathbf{q}^{j,k})$ represents the length of the $m$-th cable in the $i$-th link at configuration $\mathbf{q}^{j,k}$, and $(x)_{+}$ denotes $\max(x,0)$.
The collision cost $c_{\text{coll}}$ is defined as:
@@ VI-A Simulation Setup
To train the N-CEDF for a link of a continuum robot, we prepare the training data as described in Sec. IV-B. Each link is assumed to have an inextensible backbone length $L=2$ m and radius $r=0.2$ m, with cable length limits $l_{\text{min}}=1.6$ m and $l_{\text{max}}=2.4$ m. We uniformly sample $N=250$ configurations within the cable length limits, $N_{w}=32^{3}$ workspace points within a bounding box, and $N_{s}=1600$ surface points on the link.
@@ VII Conclusion
In this paper, we introduced a novel method for modeling the shape of continuum robots using Neural Configuration Euclidean Distance Functions (N-CEDF). By learning separate distance functions for each link and combining them through the kinematic chain, our N-CEDF efficiently and accurately represents the robot’s geometry. We integrated the N-CEDF representation with an MPPI controller for safe motion planning in dynamic and cluttered environments. Extensive simulations demonstrated the effectiveness of our approach in enabling real-time navigation of continuum robots with various numbers of links, relying solely on point cloud observations. Future work will apply the proposed method to real-world robots and explore integrating the N-CEDF representations with other planning algorithms.

### v2

@@ Abstract
This paper presents a novel method for modeling the shape of a continuum robot as a Neural Configuration Euclidean Distance Function (N-CEDF). By learning separate distance fields for each link and combining them through the kinematics chain, the learned N-CEDF provides an accurate and computationally efficient representation of the robot’s shape. The key advantage of a distance function representation of a continuum robot is that it enables efficient collision checking for motion planning in dynamic and cluttered environments, even with point-cloud observations. We integrate the N-CEDF into a Model Predictive Path Integral (MPPI) controller to generate safe trajectories for multi-segment continuum robots. The proposed approach is validated for continuum robots with various links in several simulated environments with static and dynamic obstacles.
@@ I Introduction
While many continuum robot applications focus on contact-rich interactions with the environment, this work focuses on contact-free motion planning and control, where the robot must avoid unintended contact with obstacles. Contact-free planning is critical in scenarios such as minimally invasive surgery [20, 21], where avoiding tissue damage is paramount, or in dynamic environments where the robot must proactively avoid moving obstacles [22].
We demonstrate that the advantages of the N-CEDF representation by integrating it into a Model Predictive Path Integral (MPPI) controller [23]. This enables efficient contact-free motion planning for continuum robots relying on point cloud observations of dynamic environments.
@@ III Problem Formulation
Figure 1: (a) A single continuum robot link with parameters: arc lengths $l_{1},l_{2},l_{3}$, backbone length $L$, bending angle $\theta$, and bending plane angle $\varphi$. (b) A 4-link continuum robot with a specific configuration in a dynamic environment with spherical obstacles and an end-effector goal (blue star).
A 3D multi-segment continuum robot is modeled as a series of $M$ deformable links [4, 10, 5]. Each link consists of a flexible, inextensible backbone of length $L_{i}$, actuated by three pneumatic chambers [39, 40, 41] equispaced at intervals of $\frac{2\pi}{3}$ radians around the backbone (Fig. 1(a)). The spatial configuration of each link and the entire robot body can be actively manipulated by controlling the internal pressures within these chambers. We define $l_{i,j}$, where $i\in\{1,\dots,M\}$ and $j\in\{1,2,3\}$, as the arc length along the outer surface of the continuum link, corresponding to the $j$-th chamber of link $i$. These arc lengths vary as a function of the chamber pressures due to differential expansion and contraction. We define $l_{\min}$ and $l_{\max}$ as the lower and upper bounds of these arc lengths, respectively, such that $l_{\min}\leq l_{i,j}\leq l_{\max}$ holds. With the constant curvature model, the lengths are constrained by the relationship $\frac{1}{3}\sum_{j=1}^{3}l_{i,j}=L_{i}$ for all $i$.
According to the PCC model, the bending of each link can be described by three curve parameters: the radius $\rho_{i}$ of the circular arc of the backbone, the bending angle $\theta_{i}\in[0,\pi]$, and the bending plane angle $\varphi_{i}\in[-\pi,\pi)$. These parameters can be derived from the arc lengths as follows [9, 5]:
@@
Unlike tendon-driven continuum robots, where the arc lengths $l_{i,j}$ are directly controlled through cable actuation, in pneumatic-driven robots, the arc lengths result from the internal chamber pressures $P_{i,j}$ [39, 40]. The relationship between the applied pressures and the resulting arc lengths is generally nonlinear and depends on the material properties of the robot. For simplicity, in this work, we assume that this pressure-to-arc length mapping is known or has been identified through system calibration [42]. This allows us to abstract away the underlying pressure dynamics and directly model the control inputs as changes in arc lengths.
Let $\mathbf{x}^{k}=[l_{1,1}^{k},l_{1,2}^{k},l_{1,3}^{k},\dots,l_{M,1}^{k},l_{M,2}^{k},l_{M,3}^{k}]^{\top}\in\mathbb{R}^{3M}$ denote the vector containing the arc lengths for all chambers at time step $k$. The discrete-time robot dynamics are:
@@
where $\mathbf{u}^{k}\in\mathbb{R}^{3M}$ represents the control input corresponding to changes in arc lengths, and $\tau^{k}$ is the sampling time. The relation between the configuration $\mathbf{q}^{k}$ and the arc lengths $\mathbf{x}^{k}$ is given by (1).
We consider an environment containing both static and dynamic obstacles. Let $\mathcal{O}^{k}\subset\mathbb{R}^{3}$ denote the closed obstacle set at time step $k$, and let $\mathcal{F}^{k}=\mathbb{R}^{3}\setminus\mathcal{O}^{k}$ represent the free space, which is assumed to be an open set.
@@ IV-B Data Preparation and Loss Function
To train $\hat{\Gamma}_{i}(\mathbf{p},\mathbf{q}_{i};\boldsymbol{\theta}_{i})$, we generate the following dataset. First, we uniformly sample a set of workspace points $\mathcal{P}_{w}=\{\mathbf{p}_{1}^{w},\dots,\mathbf{p}_{N_{w}}^{w}\}$ around the $i$-th link. We also uniformly sample a set of link configurations ${\cal Q}=\{\mathbf{q}_{i}^{1},\dots,\mathbf{q}_{i}^{N}\}$, where each $\mathbf{q}_{i}^{j}$ is sampled from the valid configuration space defined by the arc length limits. Next, given the configuration $\mathbf{q}_{i}^{j}$, we uniformly sample a set of points $\mathcal{P}_{s}^{j}(\mathbf{q}_{i}^{j})=\{\mathbf{p}_{1}^{s},\dots,\mathbf{p}_{N_{s}}^{s}\}$ on the surface of the link $\mathcal{B}_{i}(\mathbf{q}_{i}^{j})$. The Euclidean distance from each workspace point $\mathbf{p}_{m}^{w}\in\mathcal{P}_{w}$ to the link with configuration $\mathbf{q}_{i}^{j}$ is computed by:
@@ V-B Cost Function Design
The cost function is composed of three terms: goal-reaching cost $C_{\text{goal}}(\mathbf{Q}^{j})$, collision avoidance cost $C_{\text{coll}}(\mathbf{Q}^{j})$, and state constraint violation cost $C_{\text{state}}(\mathbf{Q}^{j})$. The goal-reaching cost penalizes the distance between the end-effector and the goal. The collision avoidance cost penalizes the robot being too close to obstacles, utilizing the learned N-CEDF model and kinematics chains. The state constraint violation cost penalizes the controlled arc lengths exceed their allowable limits. The individual terms are defined as:
@@
where $\|\cdot\|_{F}$ is the Frobenius norm, $w_{\text{goal}}$, $w_{\text{coll}}$, and $w_{\text{state}}$ are tunable weights, $l_{i,m}(\mathbf{q}^{j,k})$ represents the $m$-th arc length in the $i$-th link at configuration $\mathbf{q}^{j,k}$, and $(x)_{+}$ denotes $\max(x,0)$. The collision cost $c_{\text{coll}}$ is defined as:
@@ VI-A Simulation Setup
To train the N-CEDF for a link of a continuum robot, we prepare the training data as described in Sec. IV-B. Each link is assumed to have an inextensible backbone length $L=2$ m and radius $r=0.2$ m, with arc length limits $l_{\text{min}}=1.6$ m and $l_{\text{max}}=2.4$ m. We uniformly sample $N=250$ configurations within the arc length limits, $N_{w}=32^{3}$ workspace points within a bounding box, and $N_{s}=1600$ surface points on the link.
@@ VII Conclusion
In this paper, we introduced a novel method for modeling the shape of continuum robots using Neural Configuration Euclidean Distance Functions (N-CEDF). By learning separate distance functions for each link and combining them through the kinematic chain, our N-CEDF efficiently and accurately represents the robot’s geometry. We integrated the N-CEDF representation with an MPPI controller for safe motion planning in dynamic and cluttered environments. Extensive simulations demonstrated the effectiveness of our approach in enabling real-time navigation of continuum robots with various numbers of links, relying solely on point cloud observations. Future work will apply the proposed method to real-world continuum robots.